Hearne Company has a number of potential capital investments. Because these projects vary in nature, initial investment, and time horizon, management is finding it difficult to compare them. Assume straight line depreciation method is used.   

Project 1: Retooling Manufacturing Facility

This project would require an initial investment of $5,250,000. It would generate $937,000 in additional net cash flow each year. The new machinery has a useful life of eight years and a salvage value of $1,096,000.

Project 2: Purchase Patent for New Product

The patent would cost $3,680,000, which would be fully amortized over five years. Production of this product would generate $607,200 additional annual net income for Hearne.

Project 3: Purchase a New Fleet of Delivery Trucks

Hearne could purchase 25 new delivery trucks at a cost of $155,000 each. The fleet would have a useful life of 10 years, and each truck would have a salvage value of $5,800. Purchasing the fleet would allow Hearne to expand its customer territory resulting in $639,400 of additional net income per year.


Required:
1.
Determine each project's accounting rate of return. (Round your answers to 2 decimal places.)
projects

1

2

3


       

2. Determine each project's payback period. (Round your answers to 2 decimal places.)
projects

1

2

3


       

3. Using a discount rate of 10 percent, calculate the net present value of each project. (Future Value of $1, Present Value of $1, Future Value Annuity of $1, Present Value Annuity of $1.) (Use appropriate factor(s) from the tables provided. Round your intermediate calculations to 4 decimal places and final answers to 2 decimal places.)
Projects

1

2

3
Answer #1

Solution 1:

Accounting rate of return = Average annual income / Average investment

Average annual income:

Project 1 = $937,000 - ($5,250,000 - $1,096,000)/8 = $417,750

Project 2 = $607,200

Project 3 = $639,400

Average investment: (Cost + Salvage value)/2

Project 1 = ($5,250,000 + $1,096,000)/2 = $3,173,000

Project 2 = $3,680,000/2 = $1,840,000

Project 3 = (155000*25 + 5800*25)/2 = $2,010,000

Accounting rate of return:

Project 1 = $417,750 / $3,173,000 = 13.17%

Project 2 = $607,200 / $1,840,000 = 33%

Project 3 = $639,400 / $2,010,000 = 31.81%

Solution 2:

Payback period = Initial investment / Annual cash inflows

Annual cash inflows:

Project 1 = $937,000

Project 2 = $607,200 + ($3,680,000/5) = $1,343,200

Project 3 = $639,400 + ($3,875,000 - $145,000)/10 = $1,012,400

Payback period:

Project 1 = $5,250,000 / $937,000 = 5.60 years

Project 2 = $3,680,000 / $1,343,200 = 2.74 years

Project 3 = $3,875,000 / $1,012,400 = 3.83 years

Solution 3:

Computation of NPV - Hearne Company
Project 1 Project 2 Project 3
Particulars Period PV Factor Amount Present Value Period PV Factor Amount Present Value Period PV Factor Amount Present Value
Cash outflows:
Cost of Equipment 0 1 $5,250,000.00 $5,250,000.00 0 1 $3,680,000.00 $3,680,000.00 0 1 $3,875,000.00 $3,875,000.00
Present Value of Cash outflows (A) $5,250,000.00 $3,680,000.00 $3,875,000.00
Cash Inflows
Annual cash flows 1-8 5.3349 $937,000.00 $4,998,801.30 1-5 3.7908 $1,343,200.00 $5,091,802.56 1-10 6.1446 $1,012,400.00 $6,220,793.04
Salvage Value 8 0.4665 $1,096,000.00 $511,284.00 5 0.6209 $0.00 $0.00 10 0.3855 $145,000.00 $55,897.50
Present Value of Cash Inflows (B) $5,510,085.30 $5,091,802.56 $6,276,690.54
Net Present Value (NPV) (B-A) $260,085.30 $1,411,802.56 $2,401,690.54